Annual Meeting of the Kentucky Section of the MAA

Northern Kentucky University, Highland Heights, KY March 30-31, 2007

Contributed Talks Abstract Submission Form

All information must be received by February 28, 2007.

All contributed talks will be scheduled for 15 minutes. Please submit your name, affiliation, title, and abstract in the format illustrated in the example below. You may attach this form to an e-mail or simply type all of the information directly into an e-mail and send it to Tom Richmond at tom.richmond@wku.edu. If you have any questions, please call Tom Richmond at 270-745-6219.

An Example of an Ideal Abstract Submission:

James Barksdale, Western Kentucky University

A Plane Polygonal Law

The Pythagorean Theorem, the Law of Cosines, and the Parallelogram Law all illustrate the concept of a plane polygonal law. This presentation establishes a result which details how certain (homogeneous) plane polygonal laws create/imply other (consequential)plane figure laws.

I am:  undergrad  grad student  faculty  industry/business

E-mail:

Time Preference:  Friday afternoon  Saturday morning  None

Equipment needs:

Name, Affiliation, Title of talk, Abstract (in the format shown above):